### v1

# Learning Quantum Processes with Quantum Statistical Queries

Abstract
Learning complex quantum processes is a central challenge in many areas of quantum computing and quantum machine learning, with applications in quantum benchmarking, cryptanalysis, and variational quantum algorithms. This paper introduces the first learning framework for studying quantum process learning within the Quantum Statistical Query (QSQ) model, providing the first formal definition of statistical queries to quantum processes (QPSQs). The framework allows us to propose an efficient QPSQ learner for arbitrary quantum processes accompanied by a provable performance guarantee. We also provide numerical simulations to demonstrate the efficacy of this algorithm. The practical relevance of this framework is exemplified through application in cryptanalysis, highlighting vulnerabilities of Classical-Readout Quantum Physical Unclonable Functions (CR-QPUFs), addressing an important open question in the field of quantum hardware security. This work marks a significant step towards understanding the learnability of quantum processes and shedding light on their security implications.
1 Introduction
In recent years, the field of machine learning (ML) has experienced remarkable growth, reshaping the landscape of artificial intelligence. This transformation has left an indelible mark on natural language processing, image processing, optimization, and numerous other scientific and engineering disciplines. It has also given birth to popular and captivating tools like ChatGPT and AlphaGo. Currently, another field of research that has flourished over the past decades alongside ML is the field of Quantum computing. Quantum computing provides a fundamentally different model of computation and information processing that has given rise to creative algorithms that use these special features to provide speedups for certain computational tasks [44, 78, 2, 70], and it has unveiled a multitude of concrete and potential applications across diverse domains, including simulating complex physical systems [12], cryptanalysis [78, 17], linear optimizations [34], chemistry [57, 19] and many others. Quite naturally, the convergence of ML and quantum computing has garnered considerable attention for research in the past few years. This intersection has piqued research interest due to the presence of intriguing theoretical challenges and the practical applicability of both fields. In the realm of Quantum Machine Learning (QML), a multitude of quantum algorithms have emerged in the quest to leverage quantum machines for practical ML tasks [76, 13, 43, 83, 59, 72, 26, 4]. However, achieving a profound understanding of the foundational principles underpinning these methods and techniques remains an ongoing pursuit, perhaps not surprisingly, as the same complexity applies to both of the pillars of QML: machine learning and quantum computing. Nonetheless, there have been efforts to shine a light on the power of quantum in performing learning tasks [50, 51, 43, 45, 76, 56].
Quantum learning theory constitutes the theoretical framework to study the advantages and limitations of quantum machine learning in the broad sense of it. The main goal is to quantify learning problems as well as design and analyse algorithms to solve them. In this context of computational learning, the learner is a classical, quantum or hybrid algorithm, trying to learn an object (functions, quantum states, quantum processes, distributions, etc.) from a concept class, by having access to some examples of it. Among the different learning models in the classical learning theory literature, some have been extended to quantum and studied under the umbrella of quantum learning theory such as Quantum PAC learning [18, 8], Quantum Statistical Query model [9, 10] and Quantum agnostic learning model [8]. We refer to the following survey for an overview of the results [7]. However more broadly, this framework, in essence, encompasses a wide array of quantum learning results, ranging from learning functions encoded within quantum states [18, 11, 38], quantum state tomography [14, 66, 86], shadow tomography [48, 1] and learning diverse classes of probability distributions [24, 11, 62], to process tomography [61]. While most of the efforts in quantum learning theory have been focused on quantum states (both as examples and target objects), here in this work we shift the focus to quantum processes.
Learning quantum processes is a fundamental problem that arises in many areas in physics [84, 39] and quantum computing, such as quantum benchmarking [77, 58, 47, 15], noise characterisation [42], error mitigation [79, 71], and variational quantum algorithms [76]. Furthermore, with the crucial role of quantum computing in cryptography, another area in which the problem of learning quantum processes, particularly unitaries, arises is cryptanalysis. In these scenarios, the quantum process of interest can manifest as a quantum oracle, providing a quantum implementation of a classical function [16, 52, 75, 23, 38], or as a physical device or hardware component disclosing an unknown underlying unitary, which serves as a cryptographic key or fingerprint. [6, 68, 69]. The primary challenge in learning complex quantum processes lies in its often resource-intensive nature, rendering conventional techniques such as process tomography [61], impractical, especially for the current and near-future NISQ (Noisy Intermediate-Scale Quantum) devices. Recent endeavours have explored diverse techniques and drawn inspiration from machine learning, classical learning theory, and classical shadow methods to devise algorithms and approaches for efficiently tackling specific instances of this challenge [46, 63, 33, 77, 58, 47, 25]. However, to the best of our knowledge, no unified framework in the discipline of quantum learning theory has been established in order to formally study the learnability of quantum processes.
In this work, we provide a robust framework for the study of quantum processes in a specific learning model, namely the Quantum Statistical Query (QSQ) model, for the first time. In a statistical query model (quantum or classical) the learner constructs a hypothesis not by accessing a sequence of labelled examples themselves, but instead by adaptively querying an oracle that provides an estimate of the statistical properties of the labelled examples. In the quantum case, this statistical estimation is in fact the estimated expectation values of some observable, over multiple copies of a quantum state $\rho$, by performing efficiently implementable noisy measurements on them. This extension to the quantum world comes quite naturally as this is often the useful information extracted from a quantum system or quantum states: by measuring the quantum system several times and estimating the expectation value of an observable, which corresponds to a physical quantity. This natural correspondence to the physics of the quantum experiment and the learning tasks designed based on them marks our main motivation for the choice of this model. The feasibility of this model in practice, as compared to quantum PAC learning, makes it a good candidate for studying learning algorithms in the NISQ era and its limitations. Aside from being physically well-motivated, the model however weaker than PAC learning, is rich and interesting to study many learning problems, as it has also been a framework to show separation between quantum and classical learners. It has been shown that several concept classes are efficiently learnable in the QSQ model, whereas they require exponentially many samples with classical statistical queries [9]. Quantum statistical queries also found applications in the classical verification of quantum learning [21] and quantum error mitigation [71, 10].
Establishing our framework and definitions for learning quantum processes within the quantum statistical query model enables us to develop efficient learning algorithms for learning general quantum processes under certain distribution of states and conditions on observable. Here, our statistical query learner comes equipped with provable performance guarantees. Furthermore, we explore the practical applications of our model and learning algorithm, venturing into the domain of cryptanalysis. In this domain, our results shed light on the security vulnerabilities of a class of quantum physical unclonable functions, filling a significant gap in the study of these primitives and the protocols employing them.
1.1 Overview of the main results
We provide a list of the main contributions of this paper.
  • Framework for learning quantum processes in Quantum Statistical Query model (QPSQ): We provide a formal definition that captures what it means to learn a quantum process using statistical queries. Specifically, we define the quantum statistical query oracle QPStat for a quantum process $\mathcal{E}$ that produces a statistical estimate of the expectation value of an observable $O$, after the application of the quantum process on an input state $\rho$. We show how our oracle generalises the previously defined QStat oracles for states.
  • Efficient QPSQ learner for learning arbitrary quantum processes: We provide an algorithm that can efficiently learn arbitrary processes (under certain conditions on the observable and distribution) in the QPSQ model and hence we provide the first QPSQ learner within our framework. Our algorithm is inspired by the process-learning algorithm of Huang et al. in [46], which itself uses ideas from classical shadow tomography [48]. However, our algorithm, defined within the QPSQ model is applicable more generally and easily to a wider range of scenarios (as shown later in the application section) and provides a more physically-motivated version of the algorithm using similar classical ML methods. We also the query complexity of the algorithm as its provable performance guarantee.
  • Numerical simulations of our learning algorithm: We demonstrate the performance of our proposed algorithm through numerical simulations. We also provide methods for generating valid and simple synthetic statistical queries for our simulations.
  • Application of the QPSQ framework and learner in cryptography: We explore the applications of our model and learning algorithm in cryptanalysis. In this domain, we focus on a primitive from quantum hardware security, namely Classical-Readout of Quantum PUF or CR-QPUF. What makes this primitive a perfect case study for our model is the fact that its main security property relies on the assumption of the inherent difficulty of learning its quantum processes. This assumption has been shown to be violated for a very limited subclass of them with simple underlying unitary circuits using heuristic classical machine learning attacks [69]. However, by employing our efficient learners, we give compelling proof that CR-QPUF with even highly complex underlying quantum circuits with an exponential depth can also be learned in the QPSQ model. Notably, our results show the vulnerability of a large family of protocols relying on CR-QPUFs.
1.2 Related works
The quantum statistical query has been first introduced in [9] where they also show that parity functions, $k$-juntas and polynomial-sized DNF formulas are efficiently learnable in the QSQ model, in contrast to the classical setting where these problems are provably hard. They also introduce the notion of private PAC learning, related to differential privacy, and show its relationship to the QSQ model. The relation between quantum statistical query and differential privacy has been further investigated in [5], studying in particular local differential privacy and QSQ. Recently [10] has dived further into the QSQ model and the role of entanglement in learning in such models and providing QSQ lower bounds on learning quantum states. The paper also compares QSQ to noisy quantum PAC learning, establishing an exponential separation between the two models.
An efficient algorithm for learning the prediction of quantum processes has been introduced in [46] which uses classical shadow [48] as a subroutine. The efficiency guarantee of the algorithm is given for any quantum process under a distribution of input states that is invariant under local Clifford transformation, and with an observable that satisfies certain efficiency and locality conditions. Other approaches for learning unknown quantum processes have been suggested, such as [33], which focuses on learning quantum processes without input control, or [20], which provides methods based on the Pauli transfer matrix for learning quantum processes and Hamiltonians, emphasising the importance of quantum memory.
1.3 Organisation of the paper
We provide all the necessary background and notations that we use in the paper in Section 11. In Section 3 we introduce our quantum process learning with quantum statistical query (QPSQ) model and we provide the formal definitions and relations to other definitions. In Section 4 we provide our learning algorithm for learning arbitrary processes. We provide our numerical simulations in Section 5 and finally, the applications are discussed in Section 6.
2 Preliminaries
We start by introducing the notation we use in the paper as well as the essential background.
2.1 Quantum Information
[…]
$\mathcal{U}_{N}\mathrel{\mathop{\ordinarycolon}}=\left\{U\in\mathcal{M}_{N,N}\mathrel{\mathop{\ordinarycolon}}UU^{\dagger}=U^{\dagger}U=I\right\}.$ (1)
We will now introduce a useful orthonormal basis for $\mathcal{M}_{N,N}$ which is widely used in quantum information.
Definition 1 (Pauli operators).
The set of Pauli operators is given by
$X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$ (2)
The set $\mathcal{P}_{1}=\{I,X,Y,Z\}$ forms an orthonormal basis for $\mathcal{M}_{2,2}$ with respect to the Hilbert-Schmidt inner product. The tensor products of Pauli operators and the identity, i.e. the operators of the form $P\in\{I,X,Y,Z\}^{\otimes n}\mathrel{\mathop{\ordinarycolon}}=\mathcal{P}_{n}$, are usually referred as stabilizer operators or Pauli strings over $n$ qubits.
Definition 2 (Stabilizer states).
The eigenstates of the Pauli matrices are of special interest. We define this set as :
$stab_{1}=\{|0\rangle,|1\rangle,|+\rangle,|-\rangle,|+y\rangle,|-y\rangle\},$ (3)
where $|0\rangle$ $|1\rangle$ are the eigenstates of $Z$ , |+ket|$\rangle$| + ⟩ & |ket|$\rangle$| - ⟩ are the eigenstates of $X$ and $|+y\rangle$ $|-y\rangle$ are the eigenstates of $Y$.
Definition 3 (Clifford group).
The Clifford group is the group of unitaries generated by the following 3 gates :
$H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix},S=\begin{pmatrix}1&0\\ 0&i\end{pmatrix},CNOT=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}$ (4)
We now define the Haar measure $\mu_{N}$, which can be thought as the uniform probability distribution over all quantum states or over all unitary operators in the Hilbert space of dimension $N$. For a comprehensive introduction to the Haar measure and its properties, we refer to [60].
Definition 4 (Haar measure).
The Haar measure on the unitary group $U(N)$ is the unique probability measure $\mu_{N}$ that is both left and right invariant over the set $\mathcal{U}_{N}$, i.e., for all integrable functions $f$ and for all $V\in\mathcal{U}_{N}$, we have:
[…]
Given a state |$\ket{\phi}\in\mathbb{C}^{N}$, we denote the $k$-th moment of a Haar random state as
[…]
Note that the right invariance of the Haar measure implies that the definition of $\mathbb{E}_{\ket{\psi}\sim\mu_{N}}\left[\ket{\psi}\bra{\psi}^{\otimes k}\right]$ does not depend on the choice of |$\ket{\phi}$.
Definition 5 (Quantum process).
A map $\mathcal{E}$ from one quantum state to another is said to be completely positive if for any positive operator $A,\mathcal{E}(A)$ is also a positive operator. $\mathcal{E}$ is said to be trace-preserving if for any input density operator $\rho,tr(\mathcal{E}(\rho))=1=tr(\rho)$. A quantum process $\mathcal{E}$ is defined as a Completely Positive Trace-Preserving (CPTP) map from one quantum state to another. For a unitary $U$, the associated map is:
:$\mathcal{E}\mathrel{\mathop{\ordinarycolon}}\rho\rightarrow U\rho U^{\dagger}$ (7)
Definition 6 (POVM).
A Positive Operator-Valued Measure (POVM) is a quantum measurement described by a collection of positive operators $\{E_{m}\}_{m}$, such that $\sum_{m}E_{m}=I$ and the probability of obtaining measurement outcome $m$ on a state $\rho$ is given by $p(m)=Tr(E_{m}\rho)$.
2.2 Classical Shadow Tomography
Classical shadow tomography is the technique of using randomized measurements to learn many properties of quantum states [32, 48, 51]. It has recently been shown that classical shadow tomography can be used to predict the outcomes of arbitrary quantum processes [46]. In this section, we include relevant results on classical shadow tomography that will be used in the rest of the work.
Definition 7 (Randomized Pauli Measurement).
Given $n>0$. A randomized Pauli measurement on an $n$-qubit state is given by a $6^{n}$-outcome POVM
$\mathcal{F}^{Pauli}\triangleq\left\{\frac{1}{3^{n}}\bigotimes_{i=1}^{n}|s_{i}\rangle\langle s_{i}|\right\}_{s_{1},...,s_{n}\in stab_{1}}$ (8)
which corresponds to measuring every qubit under a random Pauli basis $(X,Y,Z$). The outcome of $\mathcal{F}^{Pauli}$ is an $n$-qubit state $|\psi\rangle=\bigotimes_{i=1}^{n}|s_{i}\rangle,$ where $|s_{i}\rangle\in stab_{1}$ is a single-qubit stabilizer state.
Next, we define classical shadows based on randomized Pauli measurements. Other measurements can also be used to define classical shadows.
Definition 8 (Classical shadow of a quantum state).
Given $n,N>0$. Consider an $n$-qubit state $\rho$. A size $N$ classical shadow of $S_{N}(\rho)$ of quantum state $\rho$ is a random set given by
$S_{N}(\rho)\triangleq\left\{|\psi_{l}\rangle\right\}_{l=1}^{N},$ (9)
where $|\psi_{l}\rangle=\bigotimes_{i=1}^{n}|s_{l,i}\rangle$ is the outcome of the $l$-th randomized Pauli measurement on a single copy of $\rho$.
Definition 9 (Classical Shadow Approximation of a quantum state).
Given the classical shadow $S_{N}(\rho)$ of an $n$-qubit state $\rho$. We can approximate $\rho$ via
$\sigma_{N}(\rho)=\frac{1}{N}\sum_{l=1}^{N}\bigotimes_{i=1}^{n}(3|s_{l,i}\rangle\langle s_{l,i}|-I)$ (10)
Definition 10 (Classical shadow of a quantum process).
Given an $n$-qubit CPTP map $\mathcal{E}$. A size-$N$ classical shadow $S_{N}(\mathcal{E})$ of the quantum process $\mathcal{E}$ is a random set given by
$S_{N}(\mathcal{E})\triangleq\left\{|\psi_{l}^{(in)}\rangle,|\psi_{l}^{(out)}\rangle\right\}_{l=1}^{N}$ (11)
where $|\psi_{l}^{(in)}\rangle=\bigotimes_{i=1}^{n}|s_{l,i}^{(in)}\rangle$ is a random input state with $|s_{l,i}^{(in)}\rangle\in stab_{1}$ sampled uniformly at random, and $|\psi_{l}^{(out)}\rangle=\bigotimes_{i=1}^{n}|s_{l,i}^{(out)}\rangle$ is the outcome of performing a random Pauli measurement on $\mathcal{E}(|\psi_{l}^{(in)}\rangle\langle\psi_{l}^{(in)}|)$.
The authors in [46] recently proposed a machine learning algorithm that is able to learn the average output behaviour of any quantum process, under some restrictions. In the learning phase, the algorithm works with the classical shadow of a generic quantum process $\mathcal{E}$ and a set of observables $\{O_{i}\}$. In the prediction phase, the algorithm receives an input quantum state $\rho$ sampled from the target distribution $D$, and aims to predict $tr(O_{i}\mathcal{E}(\rho))$ for all observables in the set. The algorithm comes with a rigorous performance guarantee on the average prediction error over $D$, achieved with efficient time and sample complexity with respect to the number of qubits and error parameters. While the guarantee holds for any quantum process, there are certain restrictions on the observables $\{O_{i}\}$ and the distribution $D$. For the detailed restrictions and the exact complexity, we refer to [46].
2.3 Computational Learning Theory
Computational learning theory studies what it means to learn a function. One of the most successful formal learning frameworks is undoubtedly the model of Probably Approximately Correct (PAC) learning, which was introduced in [82]. In this model, we consider a class of target Boolean functions $\mathcal{C}\subseteq\{f|f\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}\}$, usually called the concept class. For an arbitrary concept $c\in\mathcal{C}$, a PAC learner receives samples of the form $\{x,c(x)\}$, where, in general, $x$ is sampled from an unknown probability distribution $D\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow[0,1]$. In the setting of noisy PAC learning, the bit $c(x)$ of each sample may independently be incorrect with some probability. The learner aims to output, with high probability, a hypothesis function $h$ with low error on average over the distribution $D$.
Another learning model of interest was introduced in [53] - the Statistical Query (SQ) model. In this model, a learner is more restricted in the way it can interact with the data. Rather than learning from potentially noisy samples directly, the algorithm learns using the statistical properties of the data, making it more robust to noise. In particular, an SQ learner receives as input estimates of the expectation values of some chosen functions within specified error tolerance.
Quantum generalizations of both PAC and SQ learning have already been introduced and studied widely. The Quantum PAC (QPAC) model was introduced in [18], where the learner has access to a quantum computer and receives quantum example states as input. The quantum example state for a concept $c$ over $n$ input bits with the target distribution $D$ is the $n+1$-qubit state $|\psi_{c}\rangle=\sum_{x}\sqrt{D(x)}|x,c(x)\rangle$. It has been shown in [8] that the sample complexity of quantum and classical PAC learning is the same. However, over a fixed uniform distribution, learning with quantum queries can provide exponential advantage over the classical learner [11, 38].
A quantum analogue of statistical queries was introduced in [9]. Here, the statistical query returns an approximation of the expectation value for an input measurement observable on quantum examples of the concept class to be learned. We include the quantum statistical query oracle defined in [9] below:
Definition 11 (Qstat, from [9]).
Let $\mathcal{C}\subseteq\{c\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}\}$ be a concept class and $D\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}$ be a distribution. A quantum statistical query oracle $\mathsf{Qstat}(O,\tau)$ for some $c^{*}\in\mathcal{C}$ receives as inputs $O,\tau$, where $\tau\geq 0$ and $O\in(\mathbb{C}^{2})^{\otimes n+1}\times(\mathbb{C}^{2})^{\otimes n+1},||O||\leq 1$, and returns a number $\alpha$ satisfying
$|\alpha-\langle\psi_{c^{*}}|O|\psi_{c^{*}}\rangle|\leq\tau$
where $|\psi_{c^{*}}\rangle=\sum_{x\in\{0,1\}^{n}}\sqrt{D(x)}|x,c^{*}(x)\rangle$.
Note that in this QSQ model, while the learner can obtain an estimate of any measurement on the quantum examples, it is restricted only to classical computation. Interestingly, several concept classes such as parities, juntas, and DNF formulae are efficiently learnable in the QSQ model, whereas the classical statistical query model necessitates an exponentially larger number of samples. Additionally, the authors of [10] have established an exponential gap between QSQ learning and learning with quantum examples in the presence of classification noise.
Quantum statistical queries have also found practical applications in classical verification of quantum learning, as detailed in [21]. Furthermore, they have been employed in the analysis of quantum error mitigation models [71, 10] and quantum neural networks [31]. Alternative variations of quantum statistical queries have also been explored in [45, 37, 65].
3 Quantum Statistical Queries to Quantum Processes
In this section, we propose a framework and definition for learning quantum processes through quantum statistical queries and discuss its importance and relevance to different problems. Previously studied quantum statistical queries (Definition 11) have considered queries to quantum examples associated with some classical function [9]. While this model is the first generalisation of statistical query learning into the quantum setting and this type of query is useful for the problem of learning certain classes of classical functions, they do not encompass the quantum processes, and as such a new framework is needed for studying the learnability of quantum processes through quantum statistical queries. Learning quantum processes from (often limited amount of) data is a crucial problem in physics and many areas of quantum computing such as error characterisation and error mitigation [79, 55, 61, 42, 50]. In many realistic and near-term scenarios, the only accessible data of the quantum process is through measured outcomes of such quantum channels, which are in fact nothing but statistical queries to such quantum processes. Hence studying the quantum process learnability via statistical queries is well motivated practically from the nature of quantum experiments and measurements. In what follows, we first give a formal definition for this model and then we also clarify the relationship between our definition and the previous definition of QSQ by showing that it encompasses the latter.
Definition 12 (Statistical Query to a Quantum Process (QPSQ)).
Let :$\mathcal{E}\mathrel{\mathop{\ordinarycolon}}\mathbb{C}^{d}\rightarrow\mathbb{C}^{d}$ be a quantum process acting on a $d$-dimensional Hilbert space. A QPSQ learning algorithm has access to a quantum statistical query oracle QPStat of the process $\mathcal{E}$, which receives as input an observable $O\in\mathbb{C}^{d}\times\mathbb{C}^{d}$ satisfying $\mathinner{\!\left\lVert O\right\rVert}\leq 1$, a tolerance parameter $\tau\geq 0$, and using $poly(1/\tau)$ copies of a quantum state $\rho\in\mathcal{S}_{d}$, outputs a number $\alpha$ satisfying
[…]
We denote the query as $\alpha\leftarrow\mathsf{QPStat}_{\mathcal{E}}(\rho,O,\tau)$. The output $\alpha$ acts as an estimate of the expectation value of $O$ on the state $\rho$ after evolution under $\mathcal{E}$ within absolute error $\tau$.
Our definition of QPSQs is justified in the setting where a learner has black-box access to a quantum process, with the ability to query the process with any quantum state. We note that our definition does not specify how the copies of the target quantum state $\rho$ are being provided. The algorithm can provide multiple copies of the quantum state or can send the classical description of the quantum state to the oracle where they can be locally prepared, depending on the scenario and application. Obtaining the QPSQs for a such general class of quantum operations is then achieved by the most natural operation, which is, after evolution under this process, performing a measurement chosen by the learner. Finally, an estimate of the expectation value is returned to the learner. Again since most physical properties of a quantum system are extracted through such interactions with its associated quantum channel, the application of this model in physics is straightforward. However, we will show that this model can be used in various scenarios including quantum cryptanalysis and learning quantumly-encoded classical functions. As a practical use-case for cryptanalysis, in section 6 we demonstrate how a QPSQ learning algorithm performs as a learning attack on authentication using CR-QPUFs.
We now discuss what it means for a QPSQ learning algorithm to be efficient. Any such learning algorithm aiming to learn some property of the process $\mathcal{E}$ must be a quantum polynomial time (QPT) algorithm, making at most polynomially many queries to the oracle QPStat. As the learner itself must provide the input states or their classical description to QPStat, it must be possible to efficiently prepare the required copies of these states. This is an important point for the physical justification of this model. If the learner was able to query QPStat with arbitrary quantum states that require exponential quantum computation for preparation, the learning model would no longer be physically justified. In the following definition for an efficient QPSQ learning algorithm, we only discuss the efficiency of the algorithm, not its correctness, allowing various notions of correctness depending on the desired property of $\mathcal{E}$ to be learned up for consideration.
Definition 13 (Efficient QPSQ learner).
A QPSQ learning algorithm is called an efficient QPSQ learner if and only if it makes at most $\mathrm{poly}(\log(d))$ queries with tolerance at least $1/\mathrm{poly}(\log(d)$ to ${}_{\mathcal{E}}$ oracle, and runs in $\mathrm{poly}(\log(d))$ time.
[…]
$\textsf{QPStat}_{\mathcal{E}_{c}}(|\psi_{D}\rangle\langle\psi_{D}|,O,\tau)\equiv\textsf{Qstat}_{\ket{\psi_{c}}}(O,\tau)$ (13)
Along with the definition of Qstat, Arunachalam et. al.[9] presented algorithms for learning various concept classes in their QSQ model. The generalization of Qstat by QPStat implies that these algorithms also hold in our QPSQ learning model. Given QPStat oracle access to the process $\mathcal{E}_{c}$ associated with the target concept, and sufficient copies of the state $|\psi_{D}\rangle$, the required output from Qstat queries can be obtained using QPStat queries, and the remainder of the learning algorithms proceed identically. Thus, any concept class efficiently learnable in the QSQ model can be learned efficiently given QPSQ access to the unitary encoding of the target function instead.
4 Learning Arbitrary Processes from QPSQs
In this section, we introduce an algorithm for learning to predict the output of arbitrary quantum processes over a restricted class of observables. Our algorithm is very much inspired by the hybrid algorithm by Huang et al. in [46], which uses ideas from classical shadow tomography [48], to efficiently learn to predict the outcome of arbitrary quantum processes. In fact, our main observation was that the QPSQ model provides a more natural and general way of extracting data from a quantum process or a quantum experiment than the data extracted through the classical shadow queries, used in [46]. The classical machine learning part of our learning algorithm becomes very similar to the one in [46] with the difference being that instead of using the classical shadow approximation, we use the QPStat output to estimate the measurement outcome. This allows us to prove a rigorous performance guarantee for our learning algorithm with a modified query complexity. We discuss the applications of this learning algorithm later in section 6.
Algorithm 1 Construct statistical query database for observable $O$
for $l=1$ to $N$ do
     $|\psi^{(in)}_{l}\rangle\leftarrow\bigotimes_{i=1}^{n}|s^{(in)}_{l,i}\rangle,|s^{(in)}_{l,i}\rangle\in stab_{1}$, chosen uniformly at random
     $y_{l}(O)\leftarrow\mathsf{QPStat}_{\mathcal{E}}(|\psi^{(in)}_{l}\rangle\langle\psi^{(in)}_{l}|,O,\tau)$
end for
return $S_{N}(\mathcal{E},O)=\{|\psi^{(in)}_{l}\rangle,y_{l}(O)\}_{l=1}^{N}$
We note that the dataset we construct is specific to the observable, while the dataset shown in [46] can be reused for any valid observable. Given the classical dataset, the learning and prediction methods are identical to those from [46]. We include them here for completeness.
Algorithm 2 Learning to predict arbitrary quantum processes using statistical queries
Learning:
for all $P\in\mathcal{P}_{n},|P|\leq k=\Theta(log(1/\epsilon))$ do
     $x_{P}^{\prime\prime}(O)\leftarrow\frac{1}{N}\sum_{l=1}^{N}y_{l}(O)tr(P|\psi_{l}^{in}\rangle\langle\psi_{l}^{in}|)$
     if  $(\frac{1}{3})^{P}>2\tilde{\epsilon}$ and $|x_{P}^{\prime\prime}(O)|>2.3^{|P|/2}\sqrt{\tilde{\epsilon}}\sum_{Q\mathrel{\mathop{\ordinarycolon}}a_{Q}\neq 0}|a_{Q}|$  then
         $\hat{\alpha}_{P}(O)\leftarrow 3^{|P|}x_{P}^{\prime\prime}(O)$
     else
         $\hat{\alpha}_{P}(O)\leftarrow 0$
     end if
end for
Prediction for target state $\rho$:
$h(\rho)\leftarrow\sum_{P\mathrel{\mathop{\ordinarycolon}}|P|\leq k}\hat{\alpha}_{P}(O)tr(P\rho)$
return $h(\rho)$
We have the following rigorous performance guarantee for Algorithm 2 -
Theorem 1 (Learning an unknown quantum process using statistical queries).
Given $n,\epsilon,\tilde{\epsilon},\delta,\tau>0$ and $\tau<\tilde{\epsilon}$. Consider any unknown $n$-qubit quantum process $\mathcal{E}$ and a known $n$-qubit observable $O$ given as a sum of few-body ( $\leq\kappa=\mathcal{O}(1)$) observables, where each qubit is acted on by $\mathcal{O}(1)$ of the few-body observables. Given a training dataset $S_{N}(\mathcal{E},O)$ obtained by performing $N$ statistical queries with tolerance $\tau$ as specified in Algorithm 1, with
$N=\Omega\left(\frac{log(n^{k+\kappa}/\delta)}{(\tilde{\epsilon}-\tau)^{2}}\right)$ (14)
With probability $\geq 1-\delta$, Algorithm 2 learns a function $h(\rho)$, such that for any distribution $\mathcal{D}$ over $n$-qubit states invariant under single-qubit Clifford gates,
$\underset{\rho\sim\mathcal{D}}{\mathbb{E}}|h(\rho)-tr(O\mathcal{E}(\rho))|^{2}\leq\epsilon+\tilde{\epsilon}\max(\mathinner{\!\left\lVert O^{\prime}\right\rVert}^{2},1),$ (15)
where $O^{\prime}$ is the low-degree truncation ( of degree $k=\lceil log_{1.5}(1/\epsilon)\rceil$) of the observable $O$ after the Heisenberg evolution under $\mathcal{E}$. The training and prediction time of $h(\rho)$ are bounded above by $\mathcal{O}(kn^{k}N)$
Proof.
Our method of using statistical queries instead of classical shadow data slightly modifies the number of queries needed in the guarantee from [46]. We only show the part of the proof that varies from theirs, and refer to their work for the remainder.
Let $\mathcal{D}^{0}$ be the uniform distribution over the tensor product of $n$ single-qubit stabilizer states. Then, we define the coefficient $x_{P}(O)$ for all Paulis $P$ and observables $O$.
$x_{P}(O)=\underset{\rho\sim\mathcal{D}^{0}}{\mathbb{E}}tr(P\rho)tr(O\mathcal{E}(\rho))$ (16)
As shown in [46], these coefficients can be used to represent $O^{\prime}$, the low-degree truncation (of degree $k$) of the observable $O$ after Heisenberg evolution under $\mathcal{E}$ as defined in the theorem.
$O^{\prime}=\sum_{P\in\{I,X,Y,Z\}^{\otimes n}\mathrel{\mathop{\ordinarycolon}}|P|\leq k}3^{|P|}x_{P}(O)P$ (17)
We now begin the proof by considering a Pauli observable $Q\in\mathcal{P}_{n}$, with $|Q|\leq\kappa=\mathcal{O}(1)$. Let $\rho_{l}=|\psi_{l}^{in}\rangle\langle\psi_{l}^{in}|$. We define the random variables $x_{P}^{\prime}(Q)$ and $x_{P}^{\prime\prime}(Q)$ below:
$x_{P}^{\prime}(Q)=\frac{1}{N}\sum_{l=1}^{N}tr(P\rho_{l})tr(Q\mathcal{E}(\rho_{l}))$ (18)
$x_{P}^{\prime\prime}(Q)=\frac{1}{N}\sum_{l=1}^{N}tr(P\rho_{l})y_{l}(Q)$ (19)
where $y_{l}(Q)$ is the output of the statistical query $\mathsf{QPStat}_{\mathcal{E}}(\rho_{l},Q,\tau)$, as defined in Algorithm 1.
Consider the number of samples:
$N=\Omega\left(\frac{log(n^{k+\kappa}/\delta)}{(\tilde{\epsilon}-\tau)^{2}}\right)$ (20)
Since $P,Q$ are Paulis,
$tr(P\rho_{l})tr(Q\mathcal{E}(\rho_{l}))\in[-1,1]$ (21)
Also, over the choice of sampled input states,
$x_{P}(Q)=\mathbb{E}[x^{\prime}_{P}(Q)]$ (22)
For $\tau<\tilde{\epsilon}$, using Hoeffding’s inequality,
$\begin{split}Pr(|x^{\prime}_{P}(Q)-x_{P}(Q)|>\tilde{\epsilon}-\tau)&\leq 2exp\left(\frac{2N^{2}(\tilde{\epsilon}-\tau)^{2}}{\sum_{i=1}^{N}(2)^{2}}\right)\\ &=2\text{ }exp\left(\frac{-N(\tilde{\epsilon}-\tau)^{2}}{2}\right)\\ &\leq exp(-log(n^{k+\kappa}/\delta))\\ &=\frac{\delta}{n^{k+\kappa}}\end{split}$
By taking the Union bound over all $P,Q\in\mathcal{P}_{n},|P|\leq k,|Q|\leq\kappa$, we get
$\Pr[\forall P,Q\in\mathcal{P}_{n}\mathrel{\mathop{\ordinarycolon}}|x^{\prime}_{P}(Q)-x_{P}(Q)|<\tilde{\epsilon}-\tau]\geq 1-\delta$ (23)
Since $tr(P\rho_{l})\in\{-1,0,1\}$,
$|tr(P\rho_{l})tr(Q\mathcal{E}(\rho_{l}))-tr(P\rho_{l})y_{l}(Q)|\leq\tau$
By averaging over $l=1$ to $N$, we get
$|x_{P}^{\prime\prime}(Q)-x_{P}^{\prime}(Q)|\leq\tau$ (24)
Using triangle inequality, we get with probability at least $1-\delta$,
$|x^{\prime\prime}_{P}(Q)-x_{P}(Q)|\leq\tilde{\epsilon},\hskip 8.5359pt\forall P,Q\in\mathcal{P}_{n},|P|\leq k,|Q|\leq\kappa$ (25)
Analogous to our definition of $x_{P}^{\prime\prime}(Q)$, the proof in [46] begins by defining a variable $\hat{x}_{P}(Q)$. Their proof uses a bound similar to equation 25 with $\hat{x}_{P}(Q)$ instead. We refer to their work for the rest of the proof as it proceeds identically to theirs after this bound. ∎
By setting $\tau\leq C\tilde{\epsilon}$ for some constant $C<1$, we recover the same asymptotic complexity as [46]. Thus, using statistical queries with tolerance linear in one of the algorithm hyperparameters, we can perform the learning task with an efficient number of such queries and efficient computational time.
In certain practical cases, the following additional assumption about the output $\alpha$ of the oracle $\mathsf{QPStat}_{\mathcal{E}}$ may hold true:
$\mathbb{E}[\alpha]=Tr(O\mathcal{E}(\rho_{in}))$ (26)
Corollary 1.
Under the assumption on the output of QPStat shown in equation 26, the number of queries to QPStat of tolerance $\tau$ to achieve the error bound of the rigorous guarantee of Theorem 1 is given by
$N=\Omega\left(\frac{\tau^{2}log(n^{k+\kappa}/\delta)}{\tilde{\epsilon}^{2}}\right)$ (27)
Proof.
Under this assumption, we can eliminate the intermediate quantity $x_{P}^{\prime}(Q)$ in the previous proof, and directly apply Hoeffding’s inequality to obtain the bound on $|x_{P}^{\prime\prime}(Q)-x_{P}(Q)|$
Since $Tr(P\rho_{l})\in\{-1,0,1\},\forall\rho_{l}$
$tr(P\rho_{l})y_{l}(Q)\in[tr(P\rho_{l})tr(Q\mathcal{E}(\rho_{in}))-\tau,tr(P\rho_{l})tr(Q\mathcal{E}(\rho_{in}))+\tau],\forall l=1,2,...n$ (28)
Using Hoeffding’s inequality for a fixed P,Q,
$\begin{split}Pr(|x_{P}^{\prime\prime}(Q)-x_{P}(Q)|>\tilde{\epsilon})&\leq 2\text{ }exp\left(\frac{-2N^{2}\tilde{\epsilon}^{2}}{\sum_{i=1}^{N}\left(2\tau)^{2}\right)}\right)\\ &=2\text{ }exp\left(\frac{-N\tilde{\epsilon}^{2}}{2\tau^{2}}\right)\\ &\leq exp(-log(n^{k+\kappa}/\delta))\\ &=\frac{\delta}{n^{k+\kappa}}\end{split}$ (29)
By taking the Union bound over all $P,Q\in\mathcal{P}_{n},|P|\leq k,|Q|\leq\kappa$, we get
$Pr(|x_{P}^{\prime\prime}(Q)-x_{P}(Q)|>\tilde{\epsilon})\leq\delta$ (30)
Thus, we can guarantee with probability at least $1-\delta$,
$|x_{P}^{\prime\prime}(Q)-x_{P}(Q)|\leq\tilde{\epsilon},\hskip 8.5359pt\forall P,Q\in\mathcal{P}_{n},|P|\leq k,|Q|\leq\kappa$ (31)
5 Numerical Simulations
In this section, we demonstrate the performance of our learning algorithm presented in this work through simulations. Our simulation results allow us to apply our algorithms to learn processes. The code for our simulations is available in a public Github repository  [^1] https://github.com/chirag-w/qpsq-learning. Before presenting our simulation result, we give a remark on our approach for providing a good and realistic simulation of data obtained via statistical queries. In order to construct the output of the statistical query oracles in our simulations, we assume the oracle QPStat uses a method to estimate the expectation value of an observable, such as the ones shown in [48, 41, 49, 85]. In our simulation, in order to emulate the behaviour of these methods, we compute the true expectation value and output the result after adding a normally distributed error to it. The error is sampled from a normal distribution such that it is within the specified tolerance with high probability. We note that our learning model puts no assumptions on the error in the output of the queries, and in theory, this error can come from any arbitrary distribution as long as it is within the tolerance with high probability. However, we will argue that this simple method of generating the sample data already captures these scenarios well enough for the purpose of the simulation. We also compare this method with the use of classical shadows for evaluating the outcome of an observable [48] in Fig 1. We see that for the same target tolerance and success probabilities, the classical shadow method produces less error than that generated using a normal distribution. In fact, the normally distributed error achieves the exact value for the success probability, while any practical method would produce the same error or less, as it would come with a bound that may not be tight. Thus, our emulated oracle would produce greater deviations than in a practical implementation, implying that the real-life performance of the algorithm would only be similar to or better than the simulations. We use these emulated oracles for the simulation of the learning algorithms.
5.1 Learning Arbitrary Processes
In figure 2, we show the simulated performance of algorithm 2 in learning 10 haar-random unitaries over 6 qubits for a range of tolerances. We consider $O=Z\otimes I...\otimes I$, the Pauli-Z observable on the first qubit. We consider three distributions of target states, namely the uniform distributions over the computational basis states, the stabilizer product states and haar-random states. We can see from figure 2 that a lower QPSQ tolerance results in a lower prediction error for the same number of queries. We also see that the algorithm achieves similar performance when predicting the outcome on computational basis and stabilizer product states, even though the uniform distribution over the computational basis states is not locally flat and thus outside the performance guarantee. On the other hand, the distribution over haar-random states is within the guarantee, and the algorithm performs best on this distribution.
Figure 1: Comparison between simulated errors generated from a normal distribution and those generated by using classical shadow tomography to evaluate the EV of the Pauli-Z observable on random single-qubit stabilizer states after evolution under a fixed haar-random unitary. We fix a tolerance value $\tau=0.2$, and the probability of the deviation lying outside the tolerance, $\delta=0.0455$
(a) Computational basis states
(b) Stabilizer states
(c) Haar-random states
Figure 2: Average performance of the learning algorithm on 10 haar-random 6-qubit unitaries, in predicting the outcome of $Z_{1}$ on three target distributions
6 Application of QPSQ in Cryptanalysis of CR-QPUFs
In this section, we present applications of our QPSQ model in the realm of cryptanalysis. Specifically, we focus on a particular class of PUFs known as Classical Readout Quantum Physically Unclonable Functions (CR-QPUFs). Our objective is to investigate how our proposed algorithm can serve as an effective strategy for attacking their security. We have chosen this specific cryptographic primitive as our case study due to the inherent compatibility between the conceptual and formal framework of CR-QPUFs and the QPSQ framework. Our results showcase a concrete application of a particular learning model and its associated algorithms in the field of cryptography, thereby establishing a connection between the domains of quantum learning theory and quantum cryptography. In the next subsection, we will delve into our new attacks and the results and we will see that these outcomes not only address an ongoing open question regarding the security of different classes of quantum physically unclonable functions but also bridge the existing gap in the security analysis of CR-QPUFs, by demonstrating the unlikelihood of achieving the desired level of security for this specific class against efficient quantum adversaries.
Physical Unclonable Functions (PUFs) are hardware devices designed to resist cloning or replication, making them valuable for cryptographic tasks like authentication, identification, and fingerprinting [73, 22, 30, 28]. These devices have historically been realised in the classical realm using specific electrical circuits or optical materials [67, 40, 36, 80]. However, many of these implementations remain susceptible to various attacks, such as side-channel and machine-learning attacks [74, 35, 81, 54]. Given the inherent unclonability offered by quantum mechanics, the pursuit of physical unclonability in the quantum domain seems more appealing and natural. The concept of a Quantum PUF (QPUF) is formally introduced and analyzed in [6]. A QPUF is a quantum process that resists easy replication and, consequently, is challenging to learn from. Furthermore, it can be interacted with using general quantum states, yielding quantum outputs. Unlike classical PUFs, QPUFs can offer provable security guarantees. Nonetheless, developing practical hardware for a complete QPUF, and effectively using it, demands substantial resources like large-scale quantum processes and quantum memories that remain elusive in current quantum technology. As a result, a range of PUF variations necessitating different levels of quantum capability have been proposed and explored. Among these are the Classical Readout Quantum Physically Unclonable Functions (CR-QPUFs), examined in [68, 69]. CR-QPUFs represent a middle ground between fully quantum PUFs and classical PUFs. Their objective is to achieve the unlearnability characteristic of QPUFs while relying only on classical communication and storage. However, [69] introduces a classical machine-learning attack directed at the initial proposal for CR-QPUFs and a slightly more sophisticated quantum circuit of the same type. This expanded construction includes layers of Hadamard gates and single-qubit rotations, coupled with the inherent randomness stemming from noise and hardware imperfections. The experiment involving the application of basic regression algorithms to understand this construction, conducted on real IBM Q quantum machines as outlined in [69], indicates that this construction is learnable and hence does not provide any security. Nevertheless, because this construction entails a relatively simple and uncorrelated quantum circuit, it raises a significant unresolved question in this domain: Does the susceptibility of CR-QPUFs to learning attacks stem from the simplicity of the quantum process itself, or does its learnability arise from a more fundamental aspect —perhaps in the way CR-QPUFs have been characterised?
In this section, we attempt to provide an almost definitive and theoretical answer to this question, made possible through the application of our QPSQ framework and the algorithms we have devised. We apply these tools to various iterations of CR-QPUFs. Through this in-depth case study, we conclusively demonstrate that no matter the complexity or depth of the quantum circuit associated with CR-QPUFs, they remain susceptible to learning (within the relevant learning model we will elaborate upon), hence we cannot hope to achieve a secure protocol relying on them. The selection of CR-QPUFs as an ideal subject for our model comes from multiple reasons. Notably, these PUFs have been formally defined within the statistical query model, as noted in [69]. Our investigation reveals that the susceptibility of this type of PUF fundamentally arises from the fact that a large class of quantum processes can be efficiently learned in the QPSQ model. This spectrum includes processes such as Haar random unitaries which are random and extremely complicated to construct when scaling with the system size. Consequently, one cannot hope to evade the learning attacks. This finding, counterintuitive and surprising, highlights a critical fact regarding the security of the cryptographic schemes surrounding hardware assumptions, like physical unclonability: Security does not inherently rely on process complexity alone or the presumption that the quantum process cannot be fully and efficiently learnt by tomography. Instead, the crucial factors are the verification method and the nature of the data accessible to the adversary, which bears an important message for the design of secure protocols and subroutines based on hardware and physical assumptions.
Before we dive into the result, we give a brief overview of the desired security property associated with the physically unclonable function and their main application, namely authentication, and then we present a structure for the authentication protocols based on CR-QPUFs, which allows us to better demonstrate our cryptanalysis using QPSQs and show the effectiveness of our attacks.
6.1 Security property and attack model
As discussed in [6], the main cryptographic notion associated with PUFs in general, and particularly for the use-case of authentication and identification, is unforgeability. Essentially, unforgeability means that an adversary having access to an efficient-size database of input-output pairs (also called challenge-response pairs or CRPs in the context of PUF) of a specific function of interest, should not be able to reproduce the output of the function on a new input. This database in many cases is obtained via the adversary querying an oracle that realises the black-box access to the function. The unforgeability in cryptography is often defined as a formal game between a challenger and an adversary, and the forgery attacks are considered as algorithms (performed by the adversary) that win this game with non-negligible probability. Unforgeability is also extended to the quantum world generally in two directions: It can also be defined for quantum processes (instead of classical functions) [29] and for quantum adversaries having access to quantum oracles of a classical function [16, 3, 29]. As there are many levels and flavours of unforgeability that involve a quantum adversary, and the technical discussions about different formal definitions of it are outside the scope of this paper, we avoid analysing the security through formal unforgeability games and instead, we demonstrate our attacks using the security of the protocol. However, we note that the security inherently relies on this cryptographic notion which has a close connection to the notion of learnability, especially when considering quantum processes.
Nevertheless, the conceptual notion of unforgeability, together with the specific properties of the CR-QPUFs guide us to characterise our attack model. For our purpose, we consider two honest parties, a verifier $\mathcal{V}$, and a prover $\mathcal{P}$, communicating through a quantum or classical channel (depending on the challenge type). The purpose of the protocol is for the honest prover to prove its identity to the verifier with the promise that no quantum adversary can falsely identify themselves as the honest prover. A quantum adversary here is a quantum polynomial time (QPT) algorithm, that sits on the communication channel and can run arbitrary efficient quantum processes. The prover possesses a CR-QPUF device denoted as $\mathcal{C}$, which they will use for identification, and which is associated with the quantum CPTP map $\mathcal{E}$. The verifier on the other hand has a database of challenge-response pairs (CRPs) of $\mathcal{C}$, which is obtained by having direct access to $\mathcal{C}$ in the setup phase and recording the queries and their respective outputs. In this scenario, it is often considered that the device is later sent to the prover physically and from that point throughout the protocol will be in the possession of the prover. For a more realistic and conclusive attack model, in addition to having access to the communication channel, the adversary has also access to a polynomial-size database of CRPs of their own, which contains partial information on the process. Here, the adversarial models are often categorised into different classes depending on the level of access assumed for the adversary to obtain such data. A Weak Adversary has only access to a randomly selected set of CRPs. This model often reduces to the cases where the adversary can only obtain the data by recording the communications on the channel over a certain period of time, which is why this class of adversary is also referred to as network adversaries. A stronger security model is one where the adversary can have oracle access to the device, i.e. issuing their desired queries to the CR-QPUFs. This adversary is referred to as Adaptive Adversary. Although the adaptive adversarial model seems very strong, it is still realistic and often the desired security level when it comes to identification protocols and PUF-based schemes. A reason for this is that, as mentioned earlier, the device (in this case CR-QPUF) needs to be physically transferred at least once, which gives an adversary the chance to directly query and interact with it. As a result, in our study, we also consider the adaptive adversaries. In our model, a QPT adversary $\mathcal{A}$ can collect polynomially-many statistical queries, by issuing any desired state as input. This dataset is then used by the adversary to learn $\mathcal{C}$ in the QPSQ model in order to provide a verifiable outcome for a new challenge state during the challenge phase of the protocol, and hence break the security.
6.2 The general structure of a CR-QPUF-based authentication protocol
Given the adversarial model in the previous section, we now define a general authentication protocol for CR-QPUFs and discuss how their security (or soundness) is defined. Later we go over specific instances of such protocols for different choices of CR-QPUFs, distributions of the challenge states and observables. For the CR-QPUFs, we work with a definition similar to [69] based on statistical queries. However, in order to use our learning attacks, we need to adapt the definition to our framework. We will discuss how our framework easily and naturally generalises to this case. However, for now, we denote the CR-QPUFs as $\mathcal{C}$ and abstractly define it as a process, associated with the completely positive trace preserving (CPTP) map $\mathcal{E}$ over the $n-$qubit state space, able to produce statistical queries for any challenge in the challenge set, given an observable $O$ and a tolerance parameter $\tau$.
The protocol is run between a verifier $\mathcal{V}$ and a prover $\mathcal{P}$ as follows:
Protocol 3 The general structure of the CR-QPUF-based Authentication Protocol
  1. 1.
    Setup phase:
    1. (a)
      The Verifier $\mathcal{V}$ possesses a CR-QPUF $\mathcal{C}$ associated with the quantum process $\mathcal{E}$.
    2. (b)
      The Verifier $\mathcal{V}$ and the Prover $\mathcal{P}$ agree on an observable $O$ and a threshold $\tau$.
    3. (c)
      $\mathcal{V}$ builds a database $D$ of CRPs of $\mathcal{C}$ by querying $\mathsf{QPStat}_{\mathcal{E}}$ oracle on observable $O$, with threshold $\tau$
      For $i=1$ to $N$: For challenge $x_{i}$, a quantum state $\rho_{i}$ (or $\rho(x_{i})$), is sampled from a selected distribution $\mathcal{D}$ and prepared as multiple copies, issued to $\mathsf{QPStat}_{\mathcal{E}}$, and the respective statistical query response $y_{i}$ is recorded. Thus the database $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ is then constructed.
    4. (d)
      $\mathcal{V}$ physically sends $\mathcal{C}$ to the Prover $\mathcal{P}$.
      At this point $\mathcal{V}$ possesses database $D$ and $\mathcal{P}$ possesses the device $\mathcal{C}$.
  2. 2.
    Authentication phase:
    1. (a)
      $\mathcal{V}$ selects a CRP $(x_{i},y_{i})$ uniformly at random from $D$ and issues the challenge $x_{i}^{t}$ to $\mathcal{P}$.
      if $t=c$ the challenge is classical; if $t=q$ the challenge has been sent as multiple copies of the associate quantum state $\rho_{i}$.
    2. (b)
      $\mathcal{P}$ receives the challenge $x_{i}$ and proceeds as follows:
      if t=c: $\mathcal{P}$ creates $poly(1/\tau)$ copies of the state $\rho_{i}=\rho(x_{i})$.
      else if t=q: $\mathcal{P}$ proceeds to next step.
    3. (c)
      $\mathcal{P}$ obtains the output of the statistical query $y_{i}^{\prime}$ by issuing $\rho_{i}$ to $\mathcal{C}$. This step is similar to the setup phase.
    4. (d)
      $\mathcal{P}$ sends $y_{i}^{\prime}$ to $\mathcal{V}$.
    5. (e)
      $\mathcal{V}$ receives $y_{i}^{\prime}$ and verifies it. If $|y_{i}-y_{i}^{\prime}|\leq 2\tau$, the authentication is successfully passed. Otherwise, $\mathcal{V}$ aborts.
We also note that for a physical device such as QPUF or CR-QPUF, the statistical query oracle $\mathsf{QPStat}$ abstractly models a natural and physical interaction with the device, which is querying it with the given challenge and measuring the output quantum states on a desired observable and over several copies to estimate the expectation value of the observable. In other words, the oracle is the physical device itself and not a separate entity or implementation.
In the context of these authentication protocols utilizing CR-QPUF, we encounter two main factors governing the complexity of the underlying components. Firstly, there is the complexity of the channel representing CR-QPUF, which we have discussed in depth. Secondly, there is the choice of the observable. In our work, we assume the protocol is carried out using observables that are efficiently estimable, taking into account practical scenarios where both the verifier and prover can effectively measure and estimate the CR-QPUF’s outcome, regardless of the underlying circuit’s complexity. As such, and given that we aim to provide attacks with provable guarantees, we assume that the observable $O$, selected during the setup phase, is an efficient observable, i.e. we assume that $O$ has a polynomially bounded number of terms in its Pauli representation as well as a bounded number of qubits each term acts non-trivially on. This is a physically well-motivated assumption, as demonstrated in current state-of-the-art research on estimating the expectation value of an observable [85]. In [85], the authors provided a framework which unifies a number of the most advanced and commonly studied methods, such as those in [48, 49]. While this assumption covers a wide class of observables, there are some non-local observables that can be measured efficiently using more specific techniques, such as the one in [27]. Nevertheless, under this assumption, we are able to formally demonstrate the vulnerability of a very large class of CR-QPUF authentication protocols. However, that does not imply the security of the cases that might be excluded due to our assumption on the observable, and heuristic attacks might still be applicable to scenarios in which the protocol uses a complicated and highly non-local observable.
The correctness or completeness of the protocol, which is defined as the success probability of an honest prover in the absence of any adversary or noise over the channel, should be 1. The soundness or security of the protocol, ensures that the success probability of any adversary (depending on the adversarial model) in passing the authentication should be negligible in the security parameter. For protocols defined as above, the completeness is straightforward, hence we only define and discuss the soundness.
Definition 14 (Soundness (security) of the CR-QPUF-based Authentication Protocol).
We say the CR-QPUF-based authentication protocol defined as 3 is secure if the success probability of any QPT adaptive adversary $\mathcal{A}$ in producing an output $\tilde{y}$ for any $x$ sampled at random from a database $D$, over a distribution $\mathcal{D}$, that passes the verification i.e. satisfies the condition $|y-\tilde{y}|\leq 2\tau$, is at most negligible in the security parameter.
6.3 CR-QPUFs in QPSQ framework
While we have been discussing the QPStat oracles abstractly so far, the CR-QPUF device must, in practice, be able to take a quantum state, an observable and a tolerance parameter $\tau$, and output with high-probability a $\tau$-estimate of the expectation value. When a query is made to ${}_{\mathcal{E}}$, the device would apply $\mathcal{E}$ to multiple copies of the state, estimate the expectation value of the observable and respond with that value as output. There are multiple methods for this estimation, such as those shown in [48, 41, 49, 85], usually requiring measurements on $poly(1/\tau)$ copies of the state for statistical estimation. In a real implementation of the protocol, the device would receive copies of the input state, either through a quantum communication channel or prepared by the device given a classical description as input.
We note that there are some differences between our QPSQ model and the formal model of CR-QPUFs defined in [69]. However, we will clarify that the two models are aligned and hence our results are complimentary to their results. For this, we first recall the definition of CR-QPUF in [69]. The main difference between the two models is that in [69], the challenges have been defined in the form of descriptions of unitaries instead of quantum states. Starting with a fixed state initialized as $|0\rangle\langle 0|$, the input unitary $U_{in}$ is applied on the noisy hardware, followed by repeated measurements in the computational basis. Finally, the mean of some statistical function is computed over the measurement results. The idea behind this kind of construction is that the unique noise fingerprint of the device may result in unforgeability. We include their definition of statistical queries for CR-QPUFs below:
Definition 15 (adapted from [69]).
For a challenge unitary $U_{in}$, the response $\hat{R}_{out}$ is the mean of a set of i.i.d. samples from running $U_{in}$ on the noisy quantum device identified by $qPUF_{id}$ and measuring in the computational basis. Corresponding to the uniquely noisy device $id$ and unitary challenge $U_{in}$, we have the quantum process $\Lambda_{in}^{id}$. This gives rise to the quantum state $\rho_{in}^{id}=\Lambda_{in}^{id}(|0\rangle\langle 0|)$. Let the associated Born distribution be $\mathcal{P}_{U_{in},id}(x)=tr(M_{x}\rho_{in}^{id}M_{x}^{\dagger})$, where $M_{x}$ is the measurement operator corresponding to outcome $x$. The output of the statistical query oracle $QEval$ associated with some efficiently computable function $\phi\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}^{n}$ is then given by
$\hat{R}_{out}\leftarrow QEval(qPUF_{id},U_{in}),\text{ }|\mathbb{E}_{x\sim\mathcal{P}_{U_{in},id}}[\phi(x)]-\hat{R}_{out}|\leq\tau$ (32)
We will show that the same output distribution can be obtained from querying QPStat for an unknown noise channel $\tilde{\Lambda}^{id}$ (described below) with the input states $\rho_{in}=U_{in}|0\rangle\langle 0|U_{in}^{\dagger}$, the observable $O_{\phi}=\sum_{x}\phi(x)|x\rangle\langle x|$, and the same tolerance $\tau$.
We start by noting that Definition 15 considers a noise channel which depends on the fixed unitary input. We can model such gate-dependent noise channels as interleaved layers of unitary gates and noise. Assume the unitary $U_{in}$ is decomposed of layers of n-qubit unitary operations denoted as $U_{in}=U_{1}U_{2},\dots,U_{L}$. Then without loss of generality $\Lambda^{id}_{in}$ can be modeled as $\Lambda^{id}_{in}(\rho)=\Lambda_{L}(U_{L}\dots\Lambda_{2}(U_{2}\Lambda_{1}(U_{1}\rho U^{\dagger}_{1})U^{\dagger}_{2})\dots U^{\dagger}_{L})$. We show that is possible to rewrite this channel in the following form:
$\Lambda^{id}_{in}(\rho)=\tilde{\Lambda}_{L}\dots\tilde{\Lambda}_{2}\tilde{\Lambda}_{1}(U_{L}\dots U_{2}U_{1}\rho U^{\dagger}_{1}U^{\dagger}_{2}\dots U^{\dagger}_{L})=\tilde{\Lambda}^{id}(U_{in}\rho U^{\dagger}_{in})$ (33)
The original channel and the rearranged one have been depicted in Figure 3.
(a) Original channel
(b) Channel after rewriting
Figure 3:
In general, the intermediate noise channels ($\Lambda_{i}$) do not commute with the next unitary layers $U_{i+1}$. However, we can write each $\Lambda_{i}$ in the Kraus decomposition and carry on the non-commuting operators to the next layer, resulting in a new noise layer $\tilde{\Lambda}_{i}$ as follows:
$\Lambda_{i}(\rho)=\sum_{i}K_{i}\rho K^{\dagger}_{i},\quad\tilde{\Lambda}_{i}(\rho)=\sum_{i}\tilde{K}_{i}\rho\tilde{K}^{\dagger}_{i}$ (34)
where $K_{i}$ and $\tilde{K}_{i}$ are the Kraus operators of $\Lambda_{i}$ and $\tilde{\Lambda}_{i}$ respectively and the Kraus operators are related to each other with $\tilde{K}_{i}=U_{i+1}K_{i}U^{\dagger}_{i+1}$. Furthermore, we note that the noise channel $\Lambda^{id}_{in}$ is an unknown channel and so is the rewritten channel $\tilde{\Lambda}^{id}$. Hence, by issuing the input state $\rho_{in}^{\prime}=U_{in}|0\rangle\langle 0|$ to the oracle $\mathsf{QPStat}_{\tilde{\Lambda}^{id}}$ with observable $O_{\phi}$, we obtain a similar output to that obtained from $QEval$, showing the equivalence between the two SQ models.
$\mathsf{QPStat}_{\tilde{\Lambda}^{id}}(\rho_{in},O_{\phi},\tau)\equiv QEval(qPUF_{id},U_{in})$ (35)
6.4 Cryptanalysis results
We are now ready to analyse the security of the authentication protocol as described in Protocol 3, in our QSQ framework, with an arbitrary and potentially high-depth underlying quantum circuit for the CR-QPUF. To show the extreme case of our result, we can assume the quantum process $\mathcal{E}$ consists of a Haar-random circuit over the n-qubit Hilbert space. We can also consider any arbitrary noise model on top of it to model the hardware-specific imperfections of the CR-QPUF. However, our result holds for any arbitrary process. We use the learning algorithms 1 and 2 for our specific attack strategy. We construct the attack as follows.
Algorithm 4 Attack on Protocol 3 in QSQ with observable $O$, tolerance $\tau$
Setting hyperparameters:
$\epsilon\leftarrow\frac{\tau^{2}}{2}$
$\tilde{\epsilon}\leftarrow\frac{\tau^{2}}{2}$
Set $N$ according to Equation 14
for $i=1$ to $N$ do
     $\rho_{i}\sim stab_{1}^{\otimes n}$
     Issue challenge $\rho_{i}$ to $\mathcal{C}$
     Receive response $y_{i}$
end for
$S_{N}\leftarrow\bigl{\{}(\rho_{i},y_{i})\bigr{\}}_{i=1}^{N}$
Learn $h$ according to Algorithm 2
Forgery
Given challenge $x$ from $\mathcal{V}$, respond with $h(\rho(x))$
[…]
Theorem 2.
Protocol 3 cannot satisfy soundness, for any general choice of CR-QPUF, any choice of observable $O$ given as the sum of few-body observables, and under any choice of challenge state chosen uniformly from any locally-flat distribution (including $\mathcal{D}_{Haar}$ and $\mathcal{D}_{stab}$) as there exists an efficient QPSQ learner than produces a prediction $\tilde{y}$ for a challenge $x$, which passes the verification with non-negligible probability.
Proof.
Consider the attack described in Algorithm 4. From equation 14, we can see that the adversary makes $poly(1/\tau)$ statistical queries of tolerance $\tau$, and the algorithm runs in $poly(1/\tau)$ time. Algorithm 4 is thus an efficient QPSQ learning algorithm.
Fitting the parameters $\epsilon$ and $\tilde{\epsilon}$ into equation 15, we see that on average, for $\rho$ drawn from any locally flat distribution, and $O$ satisfying the previously stated condition with high probability,
$|h(\rho)-Tr(O\mathcal{E}(\rho))|\leq\tau$ (36)
For any query $x$ issued by $\mathcal{V}$, the associated $y$ stored in $\mathcal{D}$ was received as the output of a statistical query, implying that
$|y-Tr(O\mathcal{E}(\rho(x)))|\leq\tau$ (37)
By triangle inequality, the difference between the adversary’s prediction $h(\rho(x))$ and the stored $y$ is bounded by $2\tau$. Thus, using Algorithm 4, an adversary is able to, on average over the set of challenges, efficiently pass the authentication protocol 3 with high probability, without access to $\mathcal{C}$, breaking the soundness of the protocol. ∎
7 Conclusion and Future Works
We have formalised a quantum statistical query model for quantum processes as a physically and experimentally well-motivated learning model. This model will enable us to study the efficiency of learning algorithms, as well as the potential quantum advantages for learning problems. Our framework encompasses a natural notion of learning a quantum process, which considers all the crucial components: the process itself, the input state queries, and the observable.
One of the main advantages of our QPSQ model is that due to its generality, it allows for the formal study of various algorithms for quantum process learning which may have been designed within different disciplines and approaches. Our arbitrary process learning algorithm imposes no restrictions on the process itself but guarantees performance only for local observables and over locally flat distributions. An exciting avenue for future research is to expand this study further and seek or design algorithms with fewer restrictions on quantum process learning.
For Algorithm 2, our observations indicate that, in some applications, overcoming the locally flat distribution restriction would open up a wide range of possibilities, even if it meant restricting the family of predictable quantum processes. Conversely, in other applications, the ability to learn the process with non-local observables would be of paramount importance. The redesign of our algorithm with these alternative sets of assumptions represents an exciting area for future work, with implications for the applications of the equivalent shadow-based algorithm presented in [46].
In terms of applications, we believe there is much to explore in both proposed directions: cryptanalysis and learning theory. First, in the case of the cryptanalysis of CR-QPUFs, our attack formally breaks the security of a large class of them, for a wide range of observables. However, as discussed in section 6.2, there exists a class of efficiently estimable observables that are non-local, falling outside the formal guarantee of our attack. Therefore, while our work addresses a significant portion of the security analysis for these primitives, a narrow gap remains. It remains an open question whether it’s possible to design an efficient authentication protocol based on CR-QPUFs, using such non-local observables while maintaining resistance to learning attacks. Nonetheless, it’s worth noting that, given their definition within the statistical query model and the results presented in this paper, even if such secure protocols exist, proving their security against the powerful learning algorithms for QPSQ would be a challenging task.
Another interesting future direction is the examination of cryptographic primitives and schemes beyond CR-QPUFs through the lens of learning quantum processes with quantum statistical queries. Investigating the cryptographic implications of the efficient learners in our model as adversarial strategies presents an intriguing direction for further study.
In the realm of learning theory, our newly proposed QPSQ model provides a natural framework to study the learnability of quantum processes using statistical queries. Potential research directions using this framework include not only the study of classical SQ learning results in a quantum setting but also the adaptation of classical techniques to this setting. One such direction would be to adapt classical techniques on establishing lower bounds for learning concept classes, to learning processes in the QPSQ setting. Another compelling question in this context is whether we can show a formal separation between QPSQ and classical learners, indicating a potential path towards finding quantum advantage in learning problems.
Furthermore, it would be intriguing to observe the implementation of our learner on actual hardware or its application to data acquired from real physical experiments.
Acknowledgements: The authors thank Armando Angrisani for his valuable inputs and discussions, especially the discussions towards establishing the definition of QPSQ, and for sharing related results with us during the project. We also thank Elham Kashefi, Dominik Leichtle, Yao Ma, and Sean Thrasher for interesting discussions and comments at different stages of this project. The authors acknowledge the support of the Quantum Advantage Pathfinder (QAP), with grant reference EP/X026167/1 and the UK Engineering and Physical Sciences Research Council.